An object moves in one dimensional motion with constant acceleration a = 6.9 m/s2. At time t = 0 s, the object is at x0 = 1.7 m and has an initial velocity of v0 = 3.6 m/s.

How far will the object move before it achieves a velocity of v = 6.2 m/s?

Your answer should be accurate to the nearest 0.1 m.
Answer #1

Xo = 1.7m a= 6.9mis? Vo= 3.6m/s v= 6. 2m/s Xf- Xo using equation v2 = Vo? + 2a (X+-xo) (6.2013) = (3-6m/s)? + 2x 6.9m) 52(xp-